OpenStudy (anonymous):

Solve this: (3*n)+3=21

hartnn (hartnn):

subtract 3 from both sides, what u get ?

OpenStudy (anonymous):

How

hartnn (hartnn):

like this 3n+3-3 = 21-3

hartnn (hartnn):

so, what u get ?

OpenStudy (anonymous):

18

hartnn (hartnn):

yes, 3n =18 now divide both sides by 3

OpenStudy (anonymous):

Is it 2 am I correct

hartnn (hartnn):

no... 3n/3 =18/3 n =18/3 =... ?

OpenStudy (anonymous):

Oh I understand now it's 6

hartnn (hartnn):

yes, n=6

OpenStudy (anonymous):

Ok thank you so much

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Do u know median

hartnn (hartnn):

yes.

hartnn (hartnn):

new Q , new post please..

OpenStudy (anonymous):

Ok what is the median height in meters of 180,170,120,180,160

hartnn (hartnn):

can u arrange those numbers in ascending order ?

hartnn (hartnn):

then median is the middle value.

OpenStudy (anonymous):

Ok 120,160,170,180,180

hartnn (hartnn):

middle value is ?

OpenStudy (anonymous):

170

hartnn (hartnn):

then thats your median, 170

OpenStudy (anonymous):

A towel is 1 yard long what measure is equavilant to 1 yard in inches

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

Well, one yard is 3 feet, and 1 foot is 12 inches. so 3 x 12 is 36, 26 inches in a yard.

OpenStudy (anonymous):

Is it 36

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Zaynahf pls what is the measure of a fourth angle of a this quadrilateral 60%,110% and90%

OpenStudy (anonymous):

I think you should post it in another question

OpenStudy (anonymous):

Why

OpenStudy (anonymous):

So other can answer too, and you *kind of* have 4 questions in this post

OpenStudy (anonymous):

How do you know

OpenStudy (anonymous):

Anyone to help me

hartnn (hartnn):

in a Quadrilateral, sum of angles =360 so, let forth angle =x 60+110+90+x=360 find x from here. and please close this post.